OpenStudy (anonymous):

what is the x intercept for f(x)=x^3-x^2-x+1

OpenStudy (anonymous):

The X-Intercept is a point at which y=0. You can set your equation equal to 0 and then solve to find it. In your case, you'd go about it like this: \[y = x^3 - x^2 -x +1\] Simplifying this gives you x = 1.8393

OpenStudy (anonymous):

Remember TIP: When a + b + c + d = 0, one real root is (1). Factor y: y = x^3 - x^2 - x -1 = (x - 1)(x^2 - 1) = (x -1)^2*(x + 1). The two x-intercepts, when y = 0, have as x-coordinate: x = ..., and x -....